OpenStudy (anonymous):

Which set of information below, showing students and their letter grade in math, is NOT a function? {(Amanda, A), (Mary, B), (Florence, D), (Arianna, F)} {(Barry, B), (Thomas, B), (Barry, D), (Jane, F)} {(Dale, B), (Molly, C), (Elizabeth, A), (Doug, D)} {(Alyssa, D), (Josh, A), (Bill, D), (Olivia, F)}

OpenStudy (anonymous):

@dan815 @Joel_the_boss

JoelTheBoss (joel_the_boss):

Functions.... uhhhh lol gimme a second :P

OpenStudy (anonymous):

k

JoelTheBoss (joel_the_boss):

I think its B because the kids name is repeated twice lol

OpenStudy (anonymous):

ahh, I didn't catch that xD, thanks, I have another Q btw

JoelTheBoss (joel_the_boss):

ok tag me

OpenStudy (anonymous):

which equation is NOT a function? y = 1/3x2 1/2y = x2 − 2 x2 − 4 = y They are all functions

OpenStudy (anonymous):

@Joel_the_boss

JoelTheBoss (joel_the_boss):

@cbyer2 @One098

OpenStudy (one098):

The second one^^

JoelTheBoss (joel_the_boss):

thanks man lol Hope we helped @DUDE..LOOK..ITS..SGTC

OpenStudy (one098):

yeah

OpenStudy (one098):

..always seems that what people are taking right now I am, too, that way I can help c:

JoelTheBoss (joel_the_boss):

im learning about rationals and square rooting functions and some crazy pellet lol

OpenStudy (anonymous):

thanks